OpenStudy (anonymous):

A group of 17 people are going to an amusement park. Adult tickets cost $53.50 each and child tickets cost $47.25 each. If the total cost for admission to the park is $859.50, how many adults attended the amusement park? 7 8 9 10

OpenStudy (anonymous):

this is one expensive place but how many kids went so we can see how many adults went

OpenStudy (anonymous):

That's my problem... It doesn't say... that's the whole question. :/

OpenStudy (anonymous):

Set up two equations, one for the total number of people that went, and one for the total cost

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok i will try no promises

OpenStudy (anonymous):

@CalebBeavers What do you mean. @ashlynn143 It's okay, anything helps. (:

OpenStudy (anonymous):

Call kids x and adults y so x+y=17 since 17 people went now try to make one for the cost

OpenStudy (anonymous):

How would I make one for the cost if I don't know how many people went?

OpenStudy (anonymous):

17 people went

OpenStudy (anonymous):

Use the variables...you know the price for each one and the total cost

OpenStudy (anonymous):

could this help y=adult x= child y(53.50) + x(47.25)

OpenStudy (anonymous):

@CalebBeavers ^^^^^

OpenStudy (anonymous):

$53.50x + 47.25y = 859.50 ???

OpenStudy (anonymous):

Okay, adults can be x. thats right so now we have a system of equations x+y=17 53.50x+47.25y=859.50

OpenStudy (anonymous):

ok so now what do we do

OpenStudy (anonymous):

??? I'm still confused. I'm sorry.. I'm horrible at math and it's been a long day. X(

OpenStudy (anonymous):

because we dont know the number of kids

OpenStudy (anonymous):

I get what you have so far, but how do we get the number of adults who went?

OpenStudy (anonymous):

Solve for y in the first equation

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

How.. we don't know what x equals.

OpenStudy (anonymous):

Solve for y in terms of x, just isolate y on the left side

OpenStudy (anonymous):

okay, so... 47.25y=859.50-53.50x do I have that right so far or am I doing it wrong?

OpenStudy (anonymous):

Solve for it in the first equation x+y=17 itll be much easier

OpenStudy (anonymous):

okay so... y=17+x ... is that right so far?

OpenStudy (anonymous):

*17-x

OpenStudy (anonymous):

yep, now take 17-x and plug that in for y in the other equation

OpenStudy (anonymous):

i got to go sorry

OpenStudy (anonymous):

@ashlynn143 it's okay. (:

OpenStudy (anonymous):

What do you mean? like: 53.50x+47.25(17-x)=859.50 ???

OpenStudy (anonymous):

Yes, now solve for x

OpenStudy (anonymous):

How do I solve for x?

OpenStudy (anonymous):

Algebra...just solve like normal theres only one variable now

OpenStudy (anonymous):

yeah, but it's 17-x ... gah!! I'm confused. :(

OpenStudy (anonymous):

Do you know how to solve an equation?

OpenStudy (anonymous):

Yeah, but it's just confusing because there are no real numbers... like 17-x throws me off and gets me all confused.

OpenStudy (anonymous):

Like if I were to solve for x, I'd have to isolate the x, but I don't know how in the world to do that because the 17-x is throwing me off and I'm confused on the whole problem now. :( :(

OpenStudy (anonymous):

Distribute the 47.25 into the parentheses first

OpenStudy (anonymous):

oh... okay, so... 53.50x + 803.25 + x = 859.50 ??? right? What next? (:

OpenStudy (anonymous):

You didnt multiply the x.. 53.50x+803.25-47.25x=859.50

OpenStudy (anonymous):

ohh... okay. duh. Alright. so 53.50x + 803.25 - 47.25x = 859.50.. now what?

OpenStudy (anonymous):

combine like terms and subtract 803.25 from both sides

OpenStudy (anonymous):

okay... so I get 57.

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

thats not right

OpenStudy (anonymous):

What? :( 53.50x + 503.25 = 859.50 6.25x + 503.25= 859.50 -503.25 6.25= 356.25 +6.25 = 57

OpenStudy (anonymous):

6.25x=56.25 x=56.25/6.25

OpenStudy (anonymous):

9? ?

OpenStudy (anonymous):

(: YOU ARE AMAZING!! THANKS SO MUCH!!!! :D